OpenStudy (anonymous):

This is a real life question rather than a textbook question. Six roommates chipped in $300 each that paid a $1,800 holding fee for a house. The house however fits 8 people. So if two more roommates joined (that did not pay for the holding fee), how much would they owe towards the 6 people that already paid the holding fee?

OpenStudy (jdoe0001):

\(\bf \textit{if all 8 paid evenly the 1800}\implies \cfrac{1800}{8}=225 \\ \quad \\ \textit{now if 6 of them paid only 225}\implies 225\cdot 6=1350 \\ \quad \\ \textit{the share of the other 2 fellows would be }1800-1350\)

OpenStudy (jdoe0001):

or 225 each, since 1800/8 is 225 anyway

OpenStudy (anonymous):

The problem is, is that the 6 already paid $300 each. So I'm trying to either figure out what the other 2 roommates would owe the 6 people back.

OpenStudy (jdoe0001):

so if the other 6 already paid for the full 1800, they paid 300 each when they should have paid 225 each 300-225 = 75 so 75 * 6 = 450 = 225 * 2

OpenStudy (jdoe0001):

so the figure paid by the other 6, includes the share of the remaining two the share of the remaining two would be 225 each or 1800/8

OpenStudy (anonymous):

alright makes sense! thanks!

OpenStudy (jdoe0001):

yw